# A Survey on Inference Optimization Techniques for Mixture of Experts Models

### v1

@@ 4.2.3. Expert Loading
Expert loading time often constitutes the primary bottleneck in expert-offloading inference. To address this, some works aim to directly reduce loading costs. EdgeMoE (yi2023edgemoe, ) and HOBBIT (tang2024hobbit, ) reduce loading time by using low-precision experts instead of high-precision ones. EdgeMoE determines expert precision by profiling their importance on a dataset, while HOBBIT dynamically selects suitable precision for cache-missing experts based on current inputs. AdapMoE (zhong2024adapmoe, ) employs an adaptive gating algorithm to skip less important experts, further lowering expert loading costs.
@@ 4.2.4. CPU Assisting
Beyond relying solely on the GPU for computation, some approaches leverage the CPU to assist. Fiddler (kamahori2024fiddler, ) performs MoE computation on the CPU when the required experts are unavailable in GPU memory, instead of loading them into the GPU. MoE-Lightning (cao2024moelight, ) introduces an innovative CPU-GPU-I/O pipelining strategy, enabling simultaneous utilization of CPU and GPU resources to enhance overall system efficiency.
@@ 5. Hardware-Level Optimization
FLAME (Lin2024flame, ) is the first MoE accelerating framework which fully exploits MoE sparsity for transformer on FPGA. At the parameter level of the model, it utilizes M:N pruning to reduce unnecessary calculations, which can make balance between column balanced structured pruning and unstructured pruning; At the expert level, sparse activation prediction is performed through CEPR(circular expert prediction). By changing the patterning of activation path of experts, the accuracy of expert predictions can be effectively improved. Then use a double buffering mechanism to load the predicted expert while computing the previous expert to improve expert deployment efficiency.
M3ViT (Liang2022m3vit, ) and Edge-MoE (Sarkar2024edgemoe, ) constructed their FPGA architecture based on the reordering if attention computation in multitasking scenarios. For inference, M3ViT can activate only the sparse “expert” pathway relevant to the task of interest for efficiency, and can further achieve zero-overhead switching between tasks with their hardware-level co-design. Edge-MoE is the first end-to-end FPGA implementation for multi-task ViT proposed some aggressive techniques, including an approximate method for solving the excessive complexity of GELU function calculation on FPGA and a unified linear layer module to achieve efficient reuse of hardware resource.
Duplex (Yun2024Duplex, ) selects a destination suitable for each layer execution in the device that combines xPU and Logic PIM (processing-in-memory). that’s means it could integrate two types of processing units that share device memories. Due to the bottleneck in computation and memory access between these two processing units which can complement each other, high computation and memory access utilization can be achieved simultaneously on the same device. Besides, it introduced an alternative PIM microarchitecture. logic PIM optimized low Op/B operations through powerful processing units on the logic die, as well as more through silicon vias (TSVs) to achieve high bandwidth communication between DRAM die and logic die. Furthermore, it could excute expert and attention stage in parallel to maximize inference efficiency.

### v2

@@ 4.2.3. Expert Loading
Expert loading time often constitutes the primary bottleneck in expert-offloading inference. As reported in HOBBIT (tang2024hobbit, ), expert loading consumes more than 80% of the total inference time. To address this, several works aim to directly reduce loading costs. For instance, EdgeMoE (yi2023edgemoe, ) and HOBBIT (tang2024hobbit, ) reduce loading times by using low-precision experts instead of high-precision ones. Low-precision expert loading requires significantly less time than high-precision experts and can still maintain model accuracy, provided that only the less important experts are replaced. EdgeMoE determines the precision of each expert by profiling its importance using a static dataset. However, this method is highly dependent on the profiled dataset, which reduces its flexibility across different environments and may negatively affect accuracy. To address this limitation, HOBBIT introduces a token-level dynamic expert loading mechanism. This mechanism dynamically selects the appropriate precision for cache-missing experts based on the current inputs, ensuring both accuracy and flexibility. Specifically, it computes the importance score of a cache-missing expert based on the outputs of the gating module. If the score is below a predefined threshold, the system fetches a low-precision version; otherwise, it retrieves the high-precision version to preserve model accuracy. Additionally, AdapMoE (zhong2024adapmoe, ) employs an adaptive gating algorithm that skips less important experts. This algorithm uses the Fisher information matrix to compute the importance of each expert, further lowering expert loading costs.
@@ 4.2.4. CPU Assisting
Beyond relying solely on the GPU for computation, some approaches leverage the CPU to assist in the process. Fiddler (kamahori2024fiddler, ) performs MoE computation on the CPU by copying intermediate activation values from GPU memory to CPU memory when the required experts are unavailable in GPU memory and then transferring the outputs back to the GPU. This method eliminates the need to load missing experts into GPU memory. As measured, the latency for loading an expert from CPU memory to GPU memory is significantly higher than the combined latency of copying activation values to the CPU, performing expert computation on the CPU, and then copying the results back to the GPU. HOBBIT (tang2024hobbit, ) follows a similar pattern, using the CPU for computation with low-precision experts. Additionally, MoE-Lightning (cao2024moelight, ) introduces an innovative CPU-GPU-I/O pipelining strategy that enables the simultaneous utilization of both CPU and GPU resources, thus enhancing overall system efficiency. However, this approach is only applicable when CPU resources are sufficiently available. It may not be suitable for platforms with limited CPU resources or shared memory devices, such as the Jetson Orin.
Method Speedup Metric Baseline Main Techniques
ProMoE (song2024promoe, ) 1.16x Tokens per second LRU-MoE (song2024promoe, ) Train a predictor to prefetch experts
Fiddler (kamahori2024fiddler, ) 8.20x Tokens per second Mixtral-Offloading (eliseev2023fast, ) Use CPU to help computation
HOBBIT (tang2024hobbit, ) 2.30x Tokens per second MoE-Infinity (xue2024moe, ) Load adaptive precision experts
AdapMoE (zhong2024adapmoe, ) 1.36x Latency per sample Mixtral-Offloading (eliseev2023fast, ) Skip unimportant experts dynamically
ExpertFlow (he2024expertflow, ) 1.99x Tokens per second Cache-MoE (he2024expertflow, ) Train a predictor to prefetch experts
MoE-Infinity (xue2024moe, ) 1.96x Latency per token Mixtral-Offloading (eliseev2023fast, ) Fetch experts with request-level tracing
MoE-Lightning (cao2024moelight, ) 3.50x Tokens per second FlexGen (sheng2023flexgen, ) Schedule CPU-GPU-I/O pipeline
Mixtral-Offloading (eliseev2023fast, ) 2.28x Tokens per second Accelerate (transformers, ) Use LRU to cache important experts
Table 8. Speedup of offloading systems evaluated on Mixtral-8x7B model (jiang2024mixtral, ).
Method Speedup Metric Baseline Main Techniques
EdgeMoE (yi2023edgemoe, ) 2.82x Latency per token IO-EXP (yi2023edgemoe, ) Quantize experts into different precisions
SwapMoE (kong2023serving, ) 2.00x Latency per sample Original-MoE (kong2023serving, ) Update cached experts dynamically
SIDA (du2024sida, ) 3.93x Samples per second DeepSpeed (deepspeed, ) Prefetch experts with a hash function
ExpertFlow (he2024expertflow, ) 5.23x Tokens per second SE-MoE (shen2022se, ) Train a predictor to prefetch experts
MoE-Infinity (xue2024moe, ) 4.53x Latency per token Accelerate (transformers, ) Prefetch experts with request-level tracing
Pre-gated MoE (hwang2024pre, ) 1.50x Tokens per second MoE-OnDemand (hwang2024pre, ) Design a pre-gate MoE structure
Table 9. Speedup of offloading systems evaluated on Switch Transformer model (fedus2022switch, ).
4.2.5. Performance Summarization
Mixtral-8x7B (jiang2024mixtral, ) and the Switch Transformer (fedus2022switch, ) are popular MoE models, and most offloading systems use them as evaluation workloads. We extract relevant performance data from various papers to assess their effectiveness. Table 8 shows that many studies use Mixtral-Offloading as the baseline when evaluating the Mixtral-8x7B model. From the table, it is evident that Fiddler achieves a significant speedup compared to other systems, as it effectively utilizes the CPU to assist computation in GPU-limited environments. However, Table 9 shows that the methods when evaluated on Switch Transformer use different baselines, making direct comparisons difficult. Therefore, for our own studies on this topic, Mixtral-8x7B should stand out as a more suitable workload to consider.
4.3. Framework
Pytorch (pytorch, ) Transformers (transformers, ) DeepSpeed (deepspeed, ) Fairseq (fairseq, ) Llama.cpp (llamacpp, ) vLLM (kwon2023efficient, ) FasterTransformer (FasterTransformer, )
Parallelism System 12 5 9 2 0 0 0
Offloading System 4 7 1 0 2 1 1
Table 10. Statistics of the number of systems built on popular current frameworks.
Most of the systems mentioned above are built on top of popular deep learning frameworks such as PyTorch (pytorch, ), Transformers (transformers, ), and DeepSpeed (deepspeed, ). The data is summarized in Table 10. From the table, it is evident that most works in the expert parallelism category implement their systems based on PyTorch, DeepSpeed, and Transformers. PyTorch serves as the foundational deep learning framework, with other libraries like Transformers and DeepSpeed built on top of it. As a result, directly building a system based on PyTorch offers greater flexibility in designing custom functions, though it lacks some popular techniques found in other libraries, such as model parallelism. DeepSpeed provides numerous features to assist in model training, particularly in distributed environments. Transformers, on the other hand, offers many pretrained models, allowing developers to easily modify the model structure within the library. While this is very convenient for quickly testing ideas, it may not provide the best performance in terms of throughput and latency.
In the expert offloading category, most works still rely on Transformers due to its ease of use. However, some works also utilize Llama.cpp (llamacpp, ), vLLM (kwon2023efficient, ), and FasterTransformers (FasterTransformer, ). These three frameworks are optimized for deploying LLMs and offer great performance, but they tend to be more complex to use.
In conclusion, if you are looking to quickly verify an idea, Transformers is an excellent choice. For training models in distributed environments, DeepSpeed is the better option. If your focus is on designing an inference system with high performance, Llama.cpp and vLLM are good starting points. Otherwise, PyTorch remains a solid choice.
@@ 5. Hardware-Level Optimization
FLAME (Lin2024flame, ) is the first accelerating framework that fully exploits MoE sparsity for transformers on FPGA. At the parameter level of the model, it utilizes M:N pruning to reduce unnecessary calculations, which can make a balance between column-balanced structured pruning and unstructured pruning. At the expert level, sparse activation prediction is performed through CEPR(circular expert prediction). By changing the patterning of the activation path of experts, the accuracy of expert predictions can be effectively improved. Then use a double buffering mechanism to load the predicted expert while computing the previous expert to improve expert deployment efficiency.
M3ViT (Liang2022m3vit, ) and Edge-MoE (Sarkar2024edgemoe, ) constructed their FPGA architecture based on the reordering of attention computation in multitasking scenarios. For inference, M3ViT can activate only the sparse ”expert” pathway relevant to the task of interest for efficiency, and can further achieve zero-overhead switching between tasks with their hardware-level co-design. Edge-MoE is the first end-to-end FPGA implementation for multi-task ViT proposed some aggressive techniques, including an approximate method for solving the excessive complexity of GELU function calculation on FPGA and a unified linear layer module to achieve efficient reuse of hardware resource.
Duplex (Yun2024Duplex, ) selects a destination suitable for each layer execution in the device that combines xPU and Logic PIM (processing-in-memory). That means it could integrate two types of processing units that share device memories. Due to the bottleneck in computation and memory access between these two processing units which can complement each other, high computation and memory access utilization can be achieved simultaneously on the same device. Besides, it introduced an alternative PIM microarchitecture. The logic PIM optimized low Op/B operations through powerful processing units on the logic die, as well as more through silicon vias (TSVs) to achieve high bandwidth communication between the DRAM die and the logic die. Furthermore, it could execute expert and attention stages in parallel to maximize inference efficiency.